OpenStudy (anonymous):

Write a quadratic equation having the given numbers as a solution. -4 and -1

OpenStudy (anonymous):

(x+4)(x+1)=x^2+5x+4

OpenStudy (chaise):

If it has the solution: -4 and -1 (x+4)(x+1)=0 Foil it out. x^2+5x+4

OpenStudy (anonymous):

\[(x-(-4))(x-(-1)) = 0\]\[(x+4)(x+1) = 0\]\[x^2 + 5x + 4 = 0\]